OpenStudy (anonymous):

2x^2-x-3=0 I got x=0 x=-3/2 Did i solve correctly?

OpenStudy (anonymous):

0 cannot be the solution I think..

OpenStudy (anonymous):

Put 0 back in the equation in place of x, you will not get 0...

OpenStudy (ajprincess):

I dnt thnk x=0 is right. \(2x^2-x-3=0\) \(2x^2-3x+2x-3=0\) \(x(2x-3)+1(2x-3)=0\) \((x+1)(2x-3)=0\) \(x+1=0 ~or ~ 2x-3=0\)

OpenStudy (anonymous):

Getting @ursularoxursox

OpenStudy (anonymous):

i'm still having a difficult time understanding the equation

OpenStudy (anonymous):

i understand now. thanks for the help everyone!

OpenStudy (ajprincess):

welcome:)

OpenStudy (ajprincess):

If anythng is not clear plz ask:)

OpenStudy (anonymous):

I think I did another one wrong, can someone please guide me? The equation is 6x^2-5x-4=0 The first time I solved I got x=5/6 and x=-1 The second time I solved I got x=-1/6 and x=4 I don't think either are correct